OpenStudy (anonymous):

A laboratory experiment requires 4.8 L of a 2.5 M solution of sulfuric acid (H2SO4), but the only available H2SO4 is a 6.0 M stock solution. How could you prepare the solution needed for the lab experiment? Show all the work used to find your answer. @Abhisar

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@mayankdevnani can u plz help?

OpenStudy (mayankdevnani):

From this whole question,we can conclude that moles are same when we dilute and un-dilute acid.

OpenStudy (mayankdevnani):

correct or not ?

OpenStudy (anonymous):

yes, correct!

OpenStudy (abhisar):

\(\sf \huge M_1V_1=M_2V_2\)

OpenStudy (mayankdevnani):

so,we get \[ \bf Molarity~diluted \times volume~diluted=given~molarity \times given~volume\]

OpenStudy (anonymous):

u just copied and pasted that from yahoo answers @Nurali

OpenStudy (mayankdevnani):

understood or not? @iamabarbiegirl

OpenStudy (mayankdevnani):

just abhisar write it in short form

OpenStudy (anonymous):

oh ok

OpenStudy (abhisar):

\(\sf 4.8\times 2.5=6\times V\) Now find the value of V

OpenStudy (mayankdevnani):

then plug values,we get \[\large \bf 2.5 \times 4.8=6.0 \times x(volume)\] now solve for x

OpenStudy (abhisar):

LOL ! yes

OpenStudy (anonymous):

we divide both sidea by 6? @Abhisar

OpenStudy (mayankdevnani):

yeah and then simplify it

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\sf \huge \frac{(4.8\times 2.5)}{6}=V\) Now find the value of V \(\color{blue}{\text{End of Quote}}\) yes..

OpenStudy (anonymous):

v= 2?

OpenStudy (abhisar):

yes...

OpenStudy (abhisar):

Since we need 4.8 L. We will have to add (4.8-2=2.8 L) of water into it

OpenStudy (anonymous):

y did u subtract 2 from 4.8?

OpenStudy (abhisar):

Because we already have 2 L of 6 M solution. To make it 4.8 L of 2.5 M solution we will have to add 2.8 L of water.

OpenStudy (abhisar):

Now after adding water we will get 2+2.8=4.8 L

OpenStudy (anonymous):

ooohh okkk

OpenStudy (abhisar):

Still confused ?

OpenStudy (anonymous):

no i got it couldn't have said it better myself lool

OpenStudy (abhisar):

Sure..be honest !

OpenStudy (anonymous):

lol i am honest thats why i get all the stuff u explain to me ad am able to put in all in my own words with ur help

OpenStudy (abhisar):

that's great :) \(\huge\sf\color{green}{\text{✌゚\(\ddot\smile\) ✌゚}}\)

OpenStudy (anonymous):

haha thank you now i will put it in my own words and check it for me plz

OpenStudy (abhisar):

sure !

OpenStudy (anonymous):

does that look good? @Abhisar